You are assigned to load bar stock into a machine to make a part. Each part is 7/8" long. You have to allow an extra 1/8" for each part for the machine to cut off the part. How many feet of bar stock will you need to make 240 parts?
The machine that will make the part can only hold bars that are 4 feet long. How many bars will you have to put in the machine to make all 240 parts.
If you can only answer one of these answer the last one please.

Found 2 solutions by rfer, Theo:
Answer by rfer(16322)

A) =240"
240/12=20 ft total
20/4=5 bars

Answer by Theo(13342)

the machine will only take a 4 foot bar.

that's 48 inches per bar.

if you allow 1 inch per part (7/8 inches for the part and 1/8 inches for each cut), you can get 48 parts out of each bar.

the machine will need to make 48 cuts per bar that are 1/8 inches wide.

you will be left with 48 parts that are each 7/8 inches in length.

since you need 240 parts, and each bar will hold 48 parts, then you will need 240 / 48 = 5 bars.

the total inches of bar stock required will be 48 * 5 = 240 inches.

that comes out to be 20 feet total.

20 feet total divided by 4 equals 5 bars of 4 feet each.